Please add,putting your answer in correct scientific notation.
1.. (3.6•10^7)+(3.1•10^6)
Please multiply.
2.. (-2ab^2)(4 a^2b^2)
3.. (X+3)(x-5)
4.. -5x^2(4x-6) sorry if this is a lot and thank you

Answer by stanbon(75887)

Please add,putting your answer in correct scientific notation.
1.. (3.6•10^7)+(3.1•10^6)
= (3.6*10*10^6)+(3.1*10^6
= (36*10^6)+(3.1*10^6)
= 39.1*10*6
=============================
Please multiply.
2.. (-2ab^2)(4 a^2b^2) = -8a^3b^4
-----------------------------
3.. (X+3)(x-5) = x^2 -2x -15
-----------------------------
4.. -5x^2(4x-6) = -20x^3+30x^2
-----------------------------
Cheers,
Stan H.
---------------